The manufacturing costs of Ackerman Industries for the first three months of the year follow:

Total Costs Units Produced
January $196,560 1,690 units
February 286,320 2,690
March 305,760 4,290

Using the high-low method, determine (a) the variable cost per unit and (b) the total fixed cost. Round all answers to the nearest whole dollar.

a. Variable cost per unit $
b. Total fixed cost
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(305,760-196,560)/(4,290-1,690)

=$42 per unit

Total fixed cost=305,760-(4,290*42)=$125580

a. Variable cost per unit $42 per unit
b.Total fixed cost $125580